OpenStudy (anonymous):

v^2-2vz-24z^2 Factor trinomial

OpenStudy (amistre64):

6 and 4.. -6(4) = -24 -6+4 = -2

OpenStudy (amistre64):

(v-6z) (v+4z)

OpenStudy (amistre64):

they aint all this nice lol

OpenStudy (anonymous):

what do u mean nice (easy basically)

OpenStudy (amistre64):

yeah; these factor into nice and easy to find terms..... but they are gonna get messier as you gain confidence

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

v^2-2vz-24z^2 = 0 v^2-2vz = 24z^2 v^2-2vz +z^2 = 24z^2+z^2 (v-z)^2 = 25z^2 v-z = +-5z v = 6z, -4z (v-6z) (v+4z) is a technique that never fails ....